Find an equation of the line that passes through the given point and has the given slope. Express your answer in slope-intercept form.
(8, 0), m = -3

Answer by jim_thompson5910(35256)

y = mx + b

y = (-3)x + b

0 = (-3)*(8) + b

0 = -24 + b

0+24 = b

24 = b

b = 24

m = -3
b = 24

So the equation is y = -3x + 24